From a box containing 5 dimes and 3 nickels, 4 coins are selected at random without replacement.

Find the probability distribution for the total T of the 4 coins. Express the probability distribution graphically as a probability histogram. Find the Expected Value and Standard Deviation of the number of coins.



1
Expert's answer
2021-08-10T13:42:58-0400

For the purposes of this question, the expected probability outcomes are calculated as follows:

Let:

"D=Dimes"

"N=Nickels"

"1" "dime=10cents"

"1" "nickel=5cents"

"\u200b" "P(4D \& 0N)=\dfrac{\dbinom{5}{4}\dbinom{3}{0}}{\dbinom{5+3}{4}}=\dfrac{1}{14}"


"P(3D \& 1N)=\dfrac{\dbinom{5}{3}\dbinom{3}{1}}{\dbinom{5+3}{4}}=\dfrac{3}{7}"


"P(2D \& 2N)=\dfrac{\dbinom{5}{2}\dbinom{3}{2}}{\dbinom{5+3}{4}}=\dfrac{3}{7}"


"P(1D \& 3N)=\dfrac{\dbinom{5}{1}\dbinom{3}{3}}{\dbinom{5+3}{4}}=\dfrac{1}{14}"

The probability distribution table is presented as follows:

"E\left(T\right)=\frac{1}{14}\left(0.40\right)+\frac{3}{7}\left(0.35\right)+\frac{3}{7}\left(0.30\right)+\frac{1}{14}\left(0.25\right)"

"E\left(T\right)=0.325"


"E\left(T^2\right)=\frac{1}{14}\left(0.40\right)^2+\frac{3}{7}\left(0.35\right)^2+\frac{3}{7}\left(0.30\right)^2+\frac{1}{14}\left(0.25\right)^2"

"E\left(T^2\right)=\frac{1.7945}{14}"


"V\:ar\:T=\sigma ^2=E\left(T^2\right)-\left(E\left(T\right)\right)^2"

"V\:ar\:T=\sigma \:^2=\frac{1.4975}{14}-\left(0.325\right)^2\approx 0.001339\:\:"


"\sigma =\sqrt{\sigma ^2}=\sqrt{0.001339}\approx 0.0366"